Physics 202/208 Lab Manual

Electricity and Magnetism, Sound and Waves, Light, Modern Physics

Department of Physics University of Wisconsin, Madison

August, 2001

Copyright\begin{matrix}c \\ ◯ \end{matrix}1994 by R. Rollefson and H. Richards

ISBN 0-917974-99-9

All rights reserved. No part of this book may be reproduced, stored in a retrieval system, or transmitted in any form or by any means without permission in writing from the publisher. Printed in the USA.